13. April 2016: Mittwoch

LDAP/IdM:

  • Nightly Update: No new issues.
  • Accounts:
    • Twenty (20) errors out of 41 new accounts. Retry created them all – a reboot?
  • Issues:
    • New employee from over the weekend is confirmed. The needful was done.

Backups:

  • Save Groups:
    • Failures: MED23 (O), pespenilab (space), dpm1, pumpkinseed
    • Two of the three weekend groups running long finished overnight.

    /li>

  • Tapes:
    • Moving tapes from library to vault, vault to library
    • 194 to the Vaults
    • 200 to the Library
    • Relabeling commenced.

Tags:

Leave a Reply

You must be logged in to post a comment.